find the hypotenuse of a right triangle with sides 5 inches and 10 inches
HELP!!!

Answer by aranajay(3)

let a= 5 the first side
let b= 10 the other side
let c= ? hypotenous
from the formula
c^2 = a^2 + b^2
c^2 = 5^2 + 10^2
c^2 = 25 + 100^2
c^2 = 125-----you need to simplify
from---- 25 x 5 square root of 25 is 5
so...5 squre root(symbol)5